Upon retirement, you hope to earn a monthly income of $69,544.00. The interest rate you will earn on your investments is 7.55%, and you expect to live for 27 years after retirement. How much must you have invested at the beginning of retirement so you achieve your target income and it lasts as long as you live?

Answer =

Cash Flow =

R =

N =
Answer #1

Amount invested at the beginning of retirement is   $ 96,04,719.01

Present value of monthly income = Monthly income * Present value of annuity of 1
= $       69,544.00 * 138.109959
= $ 96,04,719.01
Working:
Present value of annuity of 1 = (1-(1+i)^-n)/i Where,
= (1-(1+0.006292)^-324)/0.006292 i = 7.55%/12 = 0.006292
= 138.109959 n = 27*12 = 324